OpenStudy (anonymous):

how much sand must be added to250lb of a cement mixture that is 22% sand to have a mixture that is 25% sand?

OpenStudy (amistre64):

250(.22) + n(1.00) = (250+n)(.25)

OpenStudy (anonymous):

okay hold on... let me see how you go about it

OpenStudy (amistre64):

250(.22) + n = 250(.25) + n(.25) n - n(.25) = 250(.25) - 250(.22) n(1 -.25) = 250(.25 - .22) n = 250 (.25 - .22) / (1 - .25)

OpenStudy (amistre64):

gotta head to class .... good luck :)

OpenStudy (anonymous):

igot10